OpenStudy (anonymous):

x +3y =6? use the x and y intercepts to graph each linear equation. please show all work.

OpenStudy (saifoo.khan):

whn x intercpets y = 0 whn y intercpets x = 0

OpenStudy (anonymous):

x+3y=6 To find the x-intercept, substitute in 0 for y and solve for x. x+3(0)=6 Multiply 3 by the 0 inside the parentheses. x+3*0=6 Multiply 3 by 0 to get 0. x+0=6 Remove the 0 from the polynomial; adding or subtracting 0 does not change the value of the expression. x=6 To find the y-intercept, substitute in 0 for x and solve for y. (0)+3y=6 Remove the parentheses around the expression 0. 0+3y=6 Adding 0 to an expression does not change the value of the expression. 3y=6 Divide each term in the equation by 3. (3y)/(3)=(6)/(3) Cancel the common factor of 3 in (3y)/(3). (<X>3<x>y)/(<X>3<x>)=(6)/(3) Remove the common factors that were cancelled out. y=(6)/(3) Cancel the common factor of 3 in (6)/(3). y=(^(2)<X>6<x>)/(<X>3<x>) Remove the common factors that were cancelled out. y=2 These are the x and y intercepts of the equation x+3y=6. x=6, y=2

OpenStudy (anonymous):

the graph is a straight line with points at 2 on the y axis and 6 on the x axis so it slants down from left to right.